A train accelerates from 0.54km/h to 0.72km/h in 10 seconds. Its acceleration is what?
1
Expert's answer
2019-11-11T15:53:09-0500

Its acceleration is 


"a=\frac{v_f-v_i}{t}"

"a=\frac{0.72-0.54}{3.6\cdot10}=0.005\frac{m}{s^2}"